Final report of the Fall 2015 semester Sabbatical Leave

Doron ZEILBERGER

During the fall Semester of 2015 I had a one-semester sabbatical leave. I visited Nankai University in Tianjin, China, and collaborated extensively with their faculty. This lead to the paper [1]. I also visited the Weiznmann Institute of Science in Israel and with my collaborators published [2].

References

[1] William Y.C. Chen, Qing-Hu Hou, and Doron Zeilberger, Automated Discovery and Proof of Congruence Theorems for Partial Sums of Combinatorial Sequences, J. of Difference Equations and Applications 22 (2016), 780-788

[2] Alon Regev, Amitai Regev, and Doron Zeilberger, Identities in character tables of S_{n}, J. of Difference Equations and its Application 22 (2016), 272-279.

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: zeilberg@math.rutgers.edu .